25mL of the 40% ethanol solution is used to prepare the 10% ethanol. What is the final volume of the solution? And, how much water must be added to do the dilution?


1
Expert's answer
2021-03-17T08:46:51-0400

Dilution factor refers to the ratio of the volume of the initial (concentrated) solution to the volume of the final (dilute) solution1, that is, the ratio of V1 to V2 or V1 : V2. A dilution factor, DF, can be calculated: DF = V2 / V1.

Dilution factor may also be expressed as the ratio of the concentration of stock solution (C1) to the concentration of the final diluted solution (C2). Stock Solution Concentration (C1) is the concentration of the stock solution.


Dillution factor is DF = V2 / V1 = C1 / C2


DF = 40 / 10 = 4


V2 = DF x V1 = 4 x 25 = 100 ml